OpenStudy (anonymous):

When an equation uses addition, the variable can be solved by the opposite process: subtraction.

OpenStudy (anonymous):

question?

OpenStudy (ybarrap):

For example x+7=0 Subtract 7 from both sides to solve x+7-7=0-7 x+0=-7 x=-7

OpenStudy (anonymous):

haley, that is correct